OpenStudy (anonymous):

How to integrate \[\int_{ - a}^a {\frac{{{x^2}}} {{\sqrt {{x^2} + {a^2}} }}dx} \]

OpenStudy (anonymous):

I have try set \[x = a\tan \theta \] but at last I have to integrate\[{a^2}\int {{{\tan }^2}\theta \sec \theta d\theta } \] which is still something I can't integrate out...

OpenStudy (anonymous):

hmm you are just making it complicated.. try to make it more simple ^.^

OpenStudy (anonymous):

Use the method of Integration by parts?

OpenStudy (anonymous):

i don't think so... but mybe it can be..

OpenStudy (anonymous):

Desperate...Take your time guys...